Calculate the maximum value of the e.m.f. generated in a coil which is rotating at 50 rev/s in a uniform magnetic field of 0.8 Wb/m2. The coil is wound on a square former having sides 5 cm in length and is wound with 300 turns


1
Expert's answer
2021-11-29T00:17:29-0500

Maximum value

"Emax= BAN\omega"

"B = 0.8 \frac{wb}{m^3}"

"A = 5\times 5 = 0.0025m^2"

"\omega = 50 \frac{rev}{s} =100 \pi \frac{rad}{s}"

"Emax= 0.08 \times 0.0025 \times 300 \times 100\pi = 188.5V"